# InvestESG: A multi-agent reinforcement learning benchmark for studying climate investment as a social dilemma

### v2

@@ 6 Discussion and Conclusion
In this paper, we introduce InvestESG, a MARL environment simulating long-term interactions between companies and investors under varying ESG disclosure policies. For policymakers, InvestESG shows that mandatory ESG disclosure, paired with highly ESG-conscious investors, can drive corporate mitigation efforts. Additionally, greenwashing poses a lesser challenge than anticipated, and providing high-quality system-wide information effectively motivates action from both corporations and investors. For economics and policy researchers, InvestESG introduces MARL as a promising tool to complement traditional empirical and theoretical methods, allowing scalable policy testing in a simulated environment. Our model predicts agent behaviors consistent with empirical evidence and uncovers novel insights. For the machine learning community, InvestESG presents a novel multi-agent benchmark, fostering the development of RL algorithms that tackle complex social dilemmas, competition, and long-term strategy—pushing forward AI applications in real-world, high-impact domains. We encourage the machine learning community to develop algorithmic innovations for InvestESG that can inspire practical actions to address climate change.
Limitations. We acknowledge that InvestESG simplifies various aspects of real-world climate evolution, financial markets, and regulatory frameworks. These simplifications were intentional for two reasons: (1) rather than fully simulating the global market and climate complexities, which is arguably infeasible for a single model, we aim to develop a "first-principles" model that offers directional insights for policy design; (2) we designed the benchmark to be simple enough for the ML community to iterate on with reasonable computational resources (e.g., a single GPU). By lowering computational barriers, we hope to encourage broader participation from the ML community.
@@ 8 Reproducibility Statement
To promote reproducibility, we provide our source code and corresponding configuration files used for running the experiments as supplementary material. The details about installation and sample code for running the experiment is also included in the supplementary material. To encourage engagement and exploration from the machine learning community, we will open source the code for the InvestESG environment and baselines via Github.

### v3

@@ 6 Discussion and Conclusion
In this paper, we introduce InvestESG, a MARL environment simulating long-term interactions between companies and investors under varying ESG disclosure policies. For policymakers, InvestESG shows that mandatory ESG disclosure, paired with highly ESG-conscious investors, can drive corporate mitigation efforts. Additionally, greenwashing poses a lesser challenge than anticipated, and providing high-quality system-wide information effectively motivates action from both corporations and investors. For economics and policy researchers, InvestESG introduces MARL as a promising tool to complement traditional empirical and theoretical methods, allowing scalable policy testing in a simulated environment. Our model predicts agent behaviors consistent with empirical evidence and uncovers novel insights. For the machine learning community, InvestESG presents a novel multi-agent benchmark, fostering the development of RL algorithms that tackle complex social dilemmas, competition, and long-term strateg, and pushing forward AI applications in real-world, high-impact domains. We encourage the machine learning community to develop algorithmic innovations for InvestESG that can inspire practical actions to address climate change.
Limitations. We acknowledge that InvestESG simplifies various aspects of real-world climate evolution, financial markets, and regulatory frameworks. These simplifications were intentional for two reasons: (1) rather than fully simulating the global market and climate complexities, which is arguably infeasible for a single model, we aim to develop a “first-principles” model that offers directional insights for policy design; (2) we designed the benchmark to be simple enough for the ML community to iterate on with reasonable computational resources (e.g., a single GPU). By lowering computational barriers, we hope to encourage broader participation from the ML community.
@@ 8 Reproducibility Statement
To promote reproducibility, we provide our source code and corresponding configuration files used for running the experiments in open source. The details about installation and sample code for running the experiment are also included at https://github.com/yuanjiayiy/InvestESG.